OpenStudy (maheshmeghwal9):

solve for {a}: [2/3(a-5)]-[1/4(a-2)]=9/2.

OpenStudy (anonymous):

Multiply throughout by 12 on both LHS and RHS and then try to solve

OpenStudy (maheshmeghwal9):

ok!

OpenStudy (anonymous):

\[2/3(a - 5) - 1/4(a - 2) = 9/2\] \[2a/3 - 10/3 - a/4 + 2 = 9/2\]

OpenStudy (anonymous):

Isolate all terms with variables on the left side and all constants on the right side.

OpenStudy (anonymous):

\[2a/3 - a/4 = 9/2 + 10/3 - 2 \]

OpenStudy (anonymous):

Do you know what to do now?

OpenStudy (maheshmeghwal9):

Will you see your first step again,please

OpenStudy (anonymous):

Did I make any mistake?

OpenStudy (maheshmeghwal9):

Ya ,u multiplied -1/4 with -2 and wrote 2

OpenStudy (anonymous):

Yes, that's right because negative multiplied by negative gives positive.

OpenStudy (anonymous):

So -1/4 * -2 = 2

OpenStudy (maheshmeghwal9):

but it should give 1/2

OpenStudy (anonymous):

Oops sorry.

OpenStudy (maheshmeghwal9):

np:(

OpenStudy (anonymous):

So its: \[2a/3 - a/4 = 9/2 + 10/3 - 1/2\]

OpenStudy (anonymous):

To clear all this fraction confusion, multiply by 12 first on both sides and then try to solve, the problem becomes much easier

OpenStudy (maheshmeghwal9):

ok 'thanx pratuo43

OpenStudy (mertsj):

\[12\times\frac{2}{3}(a-5)-12\times\frac{1}{4}(a-2)=\frac{9}{2}\times12\] \[8(a-5)-3(a-2)=54\] \[8a-40-3a+6=54\] \[5a-34=54\] \[5a=88\] \[a=\frac{88}{5}\]

OpenStudy (maheshmeghwal9):

thanx dude!

OpenStudy (mertsj):

yw